Marty is four years younger than Jennifer now in 10 years Marty will be twice the age that Jennifers now how old is each now
Answer by ikleyn(52803)

.
M = J - 4     (1)    ("Marty is four years younger than Jennifer now.")

M+10 = 2J     (2)    ("in 10 years Marty will be twice the age that Jennifers now.)


Substitute expression (1) into eq(2). You will get

J-4 + 10 = 2J  ====>  J = 6.


Then M = J-4 = 2.


Answer.  M = 2,  J = 6.